OpenStudy (anonymous):

all values of a for which the equation sin^4x+cos^4x +sin2x +a=0 is valid

OpenStudy (anonymous):

@myininaya @amistre64 @ganeshie8

OpenStudy (anonymous):

@paki

ganeshie8 (ganeshie8):

\[\begin{align}\sin^4x + \cos^4x &= (1-\cos^2x)^2 + \cos^4x \\~\\&= 2\cos^4x -2\cos^2x + 1 \\~\\&= \frac{1}{2}(2\cos^2x-1)^2 + \frac{1}{2} \\~\\&= \dfrac{\cos^2(2x)+1}{2} \end{align} \]

ganeshie8 (ganeshie8):

plug that result in the given equation, you will see a quadratic in sin(2x)

ganeshie8 (ganeshie8):

after that use below to restrict \(a\) : 1) \(\large D \ge 0\) 2) \(\large |\sin(2x)| \le 1\)

ganeshie8 (ganeshie8):

if everything goes well, you should endup with \(\large -1.5\le a\le 0.5\)